Susan wants to mix 10 pounds of Virginia nuts that cost 3.50 a pound with spanish nuts that cost 3.00 a pound to obtain a mixture that costs 3.40 a poound. How many pounds of spanish nuts should she use?
Answer by checkley75(3666)

10*3.5+3X=3.4(10+X)
35+3X=34+3.4X
3X-3.4X=34-35
-.4X=-1
X=-1/-.4
X=2.5 NUMBER OF POUNNDS OF SPANISH NUTS ARE NEEDED.
PROOF
10*3.5+3*2.5=3.4(10+2.5)
35+7.5=3.4*12.5
42.5=42.5